Determine if the statement is TRUE or FALSE. Justify your answer. All numbers under discussion are

integers.


10. Given a positive integer N ≥ 10, form the number N' by removing the ones digit from N

and subtracting this digit from the remaining truncated integer. (For example, if N = 1309,

N' = 130 0 9 = 121.) If N' is divisible by 11, then N is divisible by 11.


1
Expert's answer
2021-02-24T06:01:48-0500

We can express a positive integer "N\geq 10" as "N=10A+B" , where "B" is the ones digit.

(For example, "N=1309=10\cdot 130+9" , "A=130" and "B=9" )

Then "N^\prime =A-B" and "N^\prime" is divisible by "11" .


"N=10A+B=11A-A+B=11A-(A-B)=11A-N^\prime"

Both "11A" and "N^\prime" are divisible by "11." Therefore, "N" is divisible by "11" .